OpenStudy (anonymous):

9a-4+3(a-11)=23 (Show Work)

OpenStudy (anonymous):

9a-4+3(a-11)=23 9a-4+3*a-3*11=23 9a-4+3a-33=23 12a-37=23 12a=23+37 12a=60 a=60/12 a=5

OpenStudy (lgbasallote):

9a - 4 + 3a - 33 = 23 12a - 37 = 23 12a = 60 a = 5

OpenStudy (anonymous):

Ok, first everyone above is wrong. The answer is 2. If you substitute a for 5 it comes out 209. Just start substituting a with small numbers, like 2.

OpenStudy (lgbasallote):

9(5) - 4 + 3(5 -11) = 23 45 - 4 + 3 (-6) = 23 41 -18 =23 23 = 23 what's wrong?

OpenStudy (anonymous):

Your trying to make the problem to complicated. This is basic algebra.

OpenStudy (anonymous):

@GreenGermbubble If we sub in a=2 we get \[9(2) - 4 + 3(2 -11)\]\[ = 18 - 4 + 3 (-9) \]\[=14 -27 \]\[=-13 \neq 23\] The correct answer is 5. :)

OpenStudy (anonymous):

Haha. Your trying to make the problem more advanced than it really is. Your thinking in different terms than the problem is really in.